# Generalized Linear Bandits with Limited Adaptivity

### v3

@@ 1 Introduction
However, despite this progress, there is a key challenge that prevents deployment of CB algorithms in the real world. Practical situations often allow for very limited adaptivity, i.e., do not allow CB algorithms to update their policy at all rounds. For example, in clinical trials Group et al. (1997), each trial involves administering medical treatments to a cohort of patients, with medical outcomes observed and collected for the entire cohort at the conclusion of the trial. This data is then used to design the treatment for the next phase of the trial. Similarly, in online advertising Schwartz et al. (2017) and recommendations Li et al. (2010), updating the policy after every iteration during deployment is often infeasible due to infrastructural constraints. A recent line of work Ruan et al. (2021); Ren and Zhou (2024); Han et al. (2020); Hanna et al. (2022); Perchet et al. (2015); Gao et al. (2019); Simchi-Levi and Xu (2022) tries to address this limitation by developing algorithms that try to minimize cumulative regret while ensuring that only a limited number of policy updates occur. Across these works, two settings (called M1 and M2 from here onwards) of limited adaptivity have been popular. Both M1, M2 provide a budget $M$ to the algorithm, determining the number of times it can update its policy. In M1 Ruan et al. (2021); Hanna et al. (2023), the algorithm is required to decide upfront a sub-sequence of $M$ rounds where policy updates will occur. While in M2 Abbasi-Yadkori et al. (2011); Ruan et al. (2021)), the algorithm is allowed to adaptively decide (during its course) when to update its policy.
Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting Ruan et al. (2021); Han et al. (2020), and optimal regret guarantees were obtained when the arm feature vectors were stochastically generated. Similarly, in their seminal work on linear bandits, Abbasi-Yadkori et al. (2011) developed algorithms for the M2 setting and proved optimal regret guarantees with no restrictions on the arm vectors. While these results provide tight regret guarantees for linear reward models, extending them to generalized linear models is quite a challenge. Straightforward extensions lead to sub-optimal regret with a significantly worse dependence on an instance dependent parameter $\kappa$ (See Section 2 for definition) that captures non-linearity of the problem instance. In fact, to the best of our knowledge, developing optimal algorithms for the CB problem with generalized linear reward models under the limited adaptivity settings M1, M2, is an open research question. This is the main focus of our work. We make the following contributions.
@@ 1.1 Our Contributions
$\bullet$ We propose B-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M1 setting of limited adaptivity. We prove that, when the arm feature vectors are generated stochastically, the regret of B-GLinCB at the end of $T$ rounds is $\tilde{O}(\sqrt{T})$, when $M=\Omega(\log\log{T})$. When $M=O(\log\log{T})$, we prove an $\tilde{O}(T^{2^{M-1}/(2^{M}-2)})$ regret guarantee. While the algorithm bears a slight resemblance to the one in Ruan et al. (2021), direct utilization of their key techniques (distributional optimal design) results in a regret guarantee that scales linearly with the instance dependent non-linearity $\kappa$. On the other hand, the leading terms in our regret guarantee for B-GLinCB have no dependence on $\kappa$. To achieve this, we make novel modifications to the key technique of distributional optimal design in Ruan et al. (2021). Along with this, the rounds for policy updates are also chosen more carefully (in a $\kappa$ dependent fashion) leading to a stronger regret guarantee.
$\bullet$ We propose RS-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M2 setting of limited adaptivity. RS-GLinCB builds on a similar algorithm in Abbasi-Yadkori et al. (2011) by adding a novel context-dependent criterion for determining if a policy update is needed. This new criterion allows us to prove optimal regret guarantee ($\tilde{O}(\sqrt{T})$) with only $O(\log^{2}T)$ updates to the policy. It is quite crucial for the generalized linear reward settings, since without it the resultant regret guarantees have a linear dependence on $\kappa$.
$\bullet$ Our work also resolves a conjecture in Lee et al. (2023) by proving an optimal ($\tilde{O}(\sqrt{T})$) regret guarantee (for the CB problem with logistic reward model) that does not depend polynomially on $S$ (the known upper bound on size of the model parameters, i.e. $\|\theta^{\star}\|\leq S$, See Section 2). RS-GLinCB is, to our knowledge, the first CB algorithm for generalized linear reward models, that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret. We also perform experiments in Section 5 that validate its superiority both in terms of regret and computational efficiency in comparison to other baseline algorithms proposed in Jun et al. (2017) and Faury et al. (2020).
@@ 1.2 Important Remarks on Contributions and Comparison with Prior Work
For both B-GLinCB and RS-GLinCB, our regret guarantees are free of $\kappa$ (in their leading term), an instance-dependent parameter that can be exponential in the size of the unknown parameter vector, i.e., $\|\theta^{\star}\|$ (See Section 2 for definition). Our contribution in this regard is two-fold. Not only do we prove $\kappa$-independent regret guarantees under the limited adaptivity constraint, we also characterize a broad class of generalized linear reward models for which a $\kappa$-independent regret guarantee can be achieved. Specifically, our results imply that the CB problem with generalized linear reward models originally proposed in Filippi et al. (2010) and subsequently studied in literature (Li et al., 2017; Jun et al., 2017; Russac et al., 2020) admits a $\kappa$-independent regret.
@@
Efforts to reduce the total time complexity to be linear in $T$ have been active in the CB literature with generealized linear rewards models. For e.g., Jun et al. (2017) recently devised computationally efficient algorithms but they suffer from regret dependence on $\kappa$. Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models (Faury et al., 2020; Abeille et al., 2021), and the algorithms were subsequently made computationally efficient in Faury et al. (2022); Zhang and Sugiyama (2023). However, the techniques involved rely heavily on the structure of the logistic model and do not easily extend to more general models. To the best of our knowledge, ours is the first work that achieves optimal $\kappa$-independent regret guarantees for bounded generalized linear reward models while remaining computationally efficient [^1] While RS-GLinCB and B-GLinCB have total running time of $\widetilde{O}(T)$, their per-round complexity can reach $O(T)$. This stands in contrast to Faury et al. (2022), which maintains efficiency in both total and per-round time complexity..
@@
In order to prove $\kappa$-independent regret guarantees, we prove a key result about self concordance of bounded (almost surely) generalized linear models (Definition 2.1) in Lemma 2.2. This result was postulated in Filippi et al. (2010) for GLMs (with same definition as ours) but no proof was provided. While Faury et al. (2020; 2022) partially tackled this issue for logistic reward models [^2] In Faury (2021), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made, however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
@@ 2 Notations and Preliminaries
Important properties of GLMs such as $\operatorname*{\mathbb{E}}[r]=\dot{b}(\langle x,\theta^{\star}\rangle)$ and variance $\mathbb{V}[r]=\ddot{b}(\langle x,\theta^{\star}\rangle)$ are detailed in Appendix C. We define the link function $\mu$ as $\mu\left(\langle x,\theta^{*}\rangle\right)\coloneqq\operatorname*{\mathbb{E}}[r]$. Thus, $\mu$ is also monotone. We now present a key Lemma on GLMs (see Appendix C for details) that enable us to achieve optimal regret guarantees for our algorithms designed in Sections 3 and 4.
@@ 2.1 Instance Dependent Non-Linearity Parameters
As in prior works Faury et al. (2020; 2022), we define instance dependent parameters that capture non-linearity of the underlying instance and critically impact our algorithm design. Algorithm 1 (B-GLinCB) that solves Problem $1$, requires three such parameters that are defined using the derivative of the link function $\dot{\mu}(\cdot)$. Specifically, for any arm set ${\mathcal{X}}$, write optimal arm $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\mu\left(\langle x,\theta^{*}\rangle\right)$ and define,
$\kappa\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}({\mathcal{D}})}\max_{x\in{\mathcal{X}}}\ \frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\kappa^{*}}}\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}\left({\mathcal{D}}\right)}\ {\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\widehat{\kappa}}}\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\right]$
@@
Algorithm 2 (RS-GLinCB) that solves Problem $2$, requires another such non-linearity parameter $\kappa$, defined as,
$\kappa\coloneqq\max_{x\in\cup_{t=1}^{T}{\mathcal{X}}_{t}}\frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)}$
We note that, here, $\kappa$ is defined considering the parameter vector $\theta^{*}$ in contrast to prior work on logistic bandits Faury et al. (2022), where its definition involved a maximization over all vectors $\theta$ with $\left\lVert\theta\right\rVert\leq S$ (known upper bound of $\|\theta^{\star}\|$). Hence, our definition of $\kappa$ is potentially much smaller and can lead to lower regret, compared to prior works. Standard to the CB literature with GLM rewards, we will assume that tight upper bounds on these parameters is known to the algorithms.
@@ 2.2 Optimal Design Policies
G-optimal Design Given an arm set ${\mathcal{X}}$, the G-optimal design policy $\pi_{G}$ is the solution of the following optimization problem: $\min_{\lambda\in\Delta({\mathcal{X}})}\max_{x\in{\mathcal{X}}}{\left|\left|x\right|\right|}^{2}_{\mathbf{U}(\lambda)^{-1}}$, where $\mathbf{U}(\lambda)=\mathbb{E}_{x\sim\lambda}[xx^{\mathsf{T}}]$. Now consider the following optimization problem, also known as the D-optimal design problem: $\max_{\lambda\in\Delta({\mathcal{X}})}\log\mathsf{Det}(\mathbf{U}(\lambda))$. This is a concave maximization problem as opposed to the G-optimal design which is non-convex. We have the following equivalence theorem due to Kiefer and Wolfowitz (Kiefer and Wolfowitz, 1960):
@@
The distributional design policy is defined using a collection of tuples =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i}):p_{1},\ldots,p_{n}\geq 0\text{ and }\sum_{i}p_{i}=1\}$, wherein each ${\mathbf{M}}_{i}$ is a $d\times d$ positive semi-definite matrix and $n\leq 4d\log{d}$. The collection $\mathcal{M}$ is detailed next. Let $\text{softmax}_{\alpha}\left(\left[s_{1},\ldots,s_{k}\right]\right)$ denote the probability distribution where the $i^{th}$ element is sampled with probability $\frac{s_{i}^{\alpha}}{\sum_{j=1}^{k}s_{j}^{\alpha}}$. For a specific =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i})\}_{i=1}^{n}$, and each $i\in[n]$ write $\pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)=\text{softmax}_{\alpha}(\{\left\lVert x\right\rVert_{{\mathbf{M}}_{i}}^{2}:x\in{\mathcal{X}}\})$. Finally, with $\pi_{G}$ as the G-optimal design policy (Section 2.2), we define the Distributional optimal design policy $\pi$ as
@@
Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of (Ruan et al., 2021) to find the collection $\mathcal{M}$; see Algorithm 4 of Ruan et al. (2021). Overall, the computed $\mathcal{M}$ induces a policy $\pi$ that upholds the following guarantee.

### v4

@@ 1 Introduction
However, despite this progress, there is a key challenge that prevents deployment of CB algorithms in the real world. Practical situations often allow for very limited adaptivity, i.e., do not allow CB algorithms to update their policy at all rounds. For example, in clinical trials Group et al. (1997), each trial involves administering medical treatments to a cohort of patients, with medical outcomes observed and collected for the entire cohort at the conclusion of the trial. This data is then used to design the treatment for the next phase of the trial. Similarly, in online advertising Schwartz et al. (2017) and recommendations Li et al. (2010), updating the policy after every iteration during deployment is often infeasible due to infrastructural constraints. A recent line of work Ruan et al. (2021); Ren and Zhou (2024); Han et al. (2020); Hanna et al. (2022); Perchet et al. (2015); Gao et al. (2019); Simchi-Levi and Xu (2022) tries to address this limitation by developing algorithms that try to minimize cumulative regret while ensuring that only a limited number of policy updates occur. Across these works, two settings (called M1 and M2 from here onwards) of limited adaptivity have been popular. Both M1, M2 provide a budget $M$ to the algorithm, determining the number of times it can update its policy. In M1 Ruan et al. (2021); Hanna et al. (2023), the algorithm is required to decide upfront a sub-sequence of $M$ rounds where policy updates will occur. While in M2 Abbasi-Yadkori et al. (2011); Ruan et al. (2021)), the algorithm is allowed to adaptively decide (during its course) when to update its policy. +–+++++++++++ Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting Ruan et al. (2021); Han et al. (2020), and optimal regret guarantees were obtained when the arm feature vectors were stochastically generated. Similarly, in their seminal work on linear bandits, Abbasi-Yadkori et al. (2011) developed algorithms for the M2 setting and proved optimal regret guarantees with no restrictions on the arm vectors. While these results provide tight regret guarantees for linear reward models, extending them to generalized linear models is quite a challenge. Straightforward extensions lead to sub-optimal regret with a significantly worse dependence on an instance dependent parameter $\kappa$ (See Section 2 for definition) that captures non-linearity of the problem instance. In fact, to the best of our knowledge, developing optimal algorithms for the CB problem with generalized linear reward models under the limited adaptivity settings M1, M2, is an open research question. This is the main focus of our work. We make the following contributions.
@@ 1.1 Our Contributions
$\bullet$ We propose B-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M1 setting of limited adaptivity. We prove that, when the arm feature vectors are generated stochastically, the regret of B-GLinCB at the end of $T$ rounds is $\tilde{O}(\sqrt{T})$, when $M=\Omega(\log\log{T})$. When $M=O(\log\log{T})$, we prove an $\tilde{O}(T^{2^{M-1}/(2^{M}-2)})$ regret guarantee. While the algorithm bears a slight resemblance to the one in Ruan et al. (2021), direct utilization of their key techniques (distributional optimal design) results in a regret guarantee that scales linearly with the instance dependent non-linearity $\kappa$. On the other hand, the leading terms in our regret guarantee for B-GLinCB have no dependence on $\kappa$. To achieve this, we make novel modifications to the key technique of distributional optimal design in Ruan et al. (2021). Along with this, the rounds for policy updates are also chosen more carefully (in a $\kappa$ dependent fashion), leading to a stronger regret guarantee.
$\bullet$ We propose RS-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M2 setting of limited adaptivity. RS-GLinCB builds on a similar algorithm in Abbasi-Yadkori et al. (2011) by adding a novel context-dependent criterion for determining if a policy update is needed. This new criterion allows us to prove optimal regret guarantee ($\tilde{O}(\sqrt{T})$) with only $O(\log^{2}T)$ updates to the policy. It is quite crucial for the generalized linear reward settings since, without it, the resultant regret guarantees have a linear dependence on $\kappa$.
$\bullet$ Our work also resolves a conjecture in Lee et al. (2024) by proving an optimal ($\tilde{O}(\sqrt{T})$) regret guarantee (for the CB problem with logistic reward model) that does not depend polynomially on $S$ (the known upper bound on the size of the model parameters, i.e. $\|\theta^{\star}\|\leq S$, See Section 2111This requires a non-convex projection. We discuss its convex relaxation in Appendix E.. RS-GLinCB is, to our knowledge, the first CB algorithm for generalized linear reward models that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret. We also perform experiments in Section 5 that validate its superiority both in terms of regret and computational efficiency in comparison to other baseline algorithms proposed in Jun et al. (2017) and Faury et al. (2020).
@@ 1.2 Important Remarks on Contributions and Comparison with Prior Work
For both B-GLinCB and RS-GLinCB, our regret guarantees are free of $\kappa$ (in their leading term), an instance-dependent parameter that can be exponential in the size of the unknown parameter vector, i.e., $\|\theta^{\star}\|$ (See Section 2 for definition). Our contribution in this regard is two-fold. Not only do we prove $\kappa$-independent regret guarantees under the limited adaptivity constraint, we also characterize a broad class of generalized linear reward models for which a $\kappa$-independent regret guarantee can be achieved. Specifically, our results imply that the CB problem with generalized linear reward models originally proposed in Filippi et al. (2010) and subsequently studied in literature Li et al. (2017); Jun et al. (2017); Russac et al. (2020) admits a $\kappa$-independent regret.
@@
Efforts to reduce the total time complexity to be linear in $T$ have been active in the CB literature with generealized linear rewards models. For e.g., Jun et al. (2017) recently devised computationally efficient algorithms but they suffer from regret dependence on $\kappa$. Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models Faury et al. (2020); Abeille et al. (2021), and the algorithms were subsequently made computationally efficient in Faury et al. (2022); Zhang and Sugiyama (2023). However, the techniques involved rely heavily on the structure of the logistic model and do not easily extend to more general models. To the best of our knowledge, ours is the first work that achieves optimal $\kappa$-independent regret guarantees for bounded generalized linear reward models while remaining computationally efficient [^2] While RS-GLinCB and B-GLinCB have total running time of $\widetilde{O}(T)$, their per-round complexity can reach $O(T)$. This stands in contrast to Faury et al. (2022), which maintains efficiency in both total and per-round time complexity..
@@
In order to prove $\kappa$-independent regret guarantees, we prove a key result about self concordance of bounded (almost surely) generalized linear models (Definition 2.1) in Lemma 2.2. This result was postulated in Filippi et al. (2010) for GLMs (with the same definition as ours), but no proof was provided. While Faury et al. (2020; 2022) partially tackled this issue for logistic reward models [^3] In Faury (2021), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made; however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
@@ 2 Notations and Preliminaries
Important properties of GLMs such as $\operatorname*{\mathbb{E}}[r\mid x]=\dot{b}(\langle x,\theta^{\star}\rangle)$ and variance $\mathbb{V}[r\mid x]=\ddot{b}(\langle x,\theta^{\star}\rangle)$ are detailed in Appendix C. We define the link function $\mu$ as $\mu\left(\langle x,\theta^{*}\rangle\right)\coloneqq\operatorname*{\mathbb{E}}[r\mid x]$. Thus, $\mu$ is also monotone.
We now present a key Lemma on GLMs (see Appendix C for details) that enables us to achieve optimal regret guarantees for our algorithms designed in Sections 3 and 4.
@@ 2.1 Instance Dependent Non-Linearity Parameters
As in prior works Faury et al. (2020; 2022), we define instance dependent parameters that capture non-linearity of the underlying instance and critically impact our algorithm design. The performance of Algorithm 1 (B-GLinCB) that solves Problem $1$, can be quantified using three such parameters that are defined using the derivative of the link function $\dot{\mu}(\cdot)$. Specifically, for any arm set ${\mathcal{X}}$, write optimal arm $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\mu\left(\langle x,\theta^{*}\rangle\right)$ and define,
$\kappa\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}({\mathcal{D}})}\max_{x\in{\mathcal{X}}}\ \frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\kappa^{*}}}\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}\left({\mathcal{D}}\right)}\ {\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\widehat{\kappa}}}\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\right]$ (1)
@@
Algorithm 2 (RS-GLinCB) that solves Problem $2$, requires another such non-linearity parameter $\kappa$ [^4] We overload the notation to match that in the literature. $\kappa$ in the context of Problem 1 is defined via (1), while $\kappa$ in the context of Problem 2 by (2)., defined as,
$\kappa\coloneqq\max_{x\in\cup_{t=1}^{T}{\mathcal{X}}_{t}}\frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)}$ (2)
We note that, here, $\kappa$ is defined considering the parameter vector $\theta^{*}$ in contrast to prior work on logistic bandits Faury et al. (2022), where its definition involved a maximization over all vectors $\theta$ with $\left\lVert\theta\right\rVert\leq S$ (known upper bound of $\|\theta^{\star}\|$). Hence, $\kappa$ as defined here is potentially much smaller and can lead to lower regret, compared to prior works. Standard to the CB literature with GLM rewards, we will assume that tight upper bounds on these parameters is known to the algorithms.
@@ 2.2 Optimal Design Policies
G-optimal Design Given an arm set ${\mathcal{X}}$, the G-optimal design policy $\pi_{G}$ is the solution of the following optimization problem: $\operatorname*{arg\,min}_{\lambda\in\Delta({\mathcal{X}})}\max_{x\in{\mathcal{X}}}{\left|\left|x\right|\right|}^{2}_{\mathbf{U}(\lambda)^{-1}}$, where $\mathbf{U}(\lambda)=\mathbb{E}_{x\sim\lambda}[xx^{\mathsf{T}}]$. Now consider the following optimization problem, also known as the D-optimal design problem: $\max_{\lambda\in\Delta({\mathcal{X}})}\log\mathsf{Det}(\mathbf{U}(\lambda))$. This is a concave maximization problem as opposed to the G-optimal design which is non-convex. We have the following equivalence theorem due to Kiefer and Wolfowitz Kiefer and Wolfowitz (1960):
@@
The distributional optimal design policy is defined using a collection of tuples =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i}):p_{1},\ldots,p_{n}\geq 0\text{ and }\sum_{i}p_{i}=1\}$, wherein each ${\mathbf{M}}_{i}$ is a $d\times d$ positive semi-definite matrix and $n\leq 4d\log{d}$. The collection $\mathcal{M}$ is detailed next. Let $\text{softmax}_{\alpha}\left(\left\{s_{1},\ldots,s_{k}\right\}\right)$ denote the probability distribution where the $i^{th}$ element is sampled with probability $\frac{s_{i}^{\alpha}}{\sum_{j=1}^{k}s_{j}^{\alpha}}$. For a specific =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i})\}_{i=1}^{n}$, and each $i\in[n]$ write $\pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)=\text{softmax}_{\alpha}(\{\left\lVert x\right\rVert_{{\mathbf{M}}_{i}}^{2}:x\in{\mathcal{X}}\})$. Finally, with $\pi_{G}$ as the G-optimal design policy (Section 2.2), we define the Distributional optimal design policy $\pi$ as
@@
Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of Ruan et al. (2021) to find the collection $\mathcal{M}$; see Algorithm 4 of Ruan et al. (2021). Overall, the computed $\mathcal{M}$ induces a policy $\pi$ that upholds the following guarantee.